A COCYCLE FORMULA FOR THE QUATERNIONIC DISCRETE SERIES

ROBERT W. DONLEY JR.

Abstract. Schmid’s proof of the Kostant-Langlands conjecture for discrete series representations of a semisimple Lie group provides a Hilbert space realization for such representations in L^{2}-cohomology. We give an explicit description of these harmonic forms for the quaternionic discrete series.